SLaTtvAmP96:

A white-tailed deer can sprint at speeds up to 30 miles per hour. American bison can run at speeds up to 3,520 feet per minute. Which animal is faster and by how many miles per hour? There are 5,280 feet in one mile.

NerdyLizard:

You have to convert the values to be the same. 60 minutes in an hour. and 5280 feet in a mile. Good luck! :)

crispyrat:

following on to what they said \(3,520*60\) is how much a bison can do in an hour so you would get \(211200\) and since there is \(5,280\) in an feet then that means a deer can go \(5,280*30=158400\) so the nison goes faster by \(211200-158400=52800\)

mhanifa:

We need both speeds in miles per hour. Bison's speed is 3520 feet per minute. This needs converted. We know 1 mile = 5280 feet, 1 hour = 60 min. Then 1 foot / min = 1/5280 mile : 1/60 hour >> 3520 feet/min = 60*3520/5280 = 40 miles/hour Now we can compare: 40 - 30 = 10 American bison is faster by 10 miles per hour.